# Thermal nature of confining strings

### v1

@@ I Introduction
Hadron production in high-energy collisions exhibits thermal features that appear universal and do not depend on the nature of colliding particles and the collision energy. The yields of different hadrons produced in $e^{+}e^{-}$, proton-proton, and heavy-ion collisions are well described by statistical models using a universal temperature Becattini (1997); Andronic et al. (2009). This apparent thermal behavior is very unlikely to originate from a conventional thermalization process involving rescattering of produced particles. For example, in $e^{+}e^{-}$ annihilation into hadrons, the produced hadrons (and the preceding quarks and gluons) are separated by a space-like interval and cannot undergo multiple final-state interactions.
Moreover, strong elliptical flow observed in heavy ion collisions at RHIC and LHC points towards a very fast thermalization – otherwise, the hydrodynamical evolution is too short and is not capable to reproduce the data, see Schlichting and Teaney (2019) for a review. This early thermalization challenges the conventional models based on kinetic theory Schlichting and Teaney (2019).
@@
Concrete realizations of this idea include the entanglement produced a rapid pulse of (chromo-)electric field during a high-energy collision Kharzeev and Tuchin (2005); Kharzeev et al. (2007); Castorina et al. (2007); Florio and Kharzeev (2021); Grieninger et al. (2023a, b). In this case an effective temperature $T$ appears proportional to the inverse duration of the pulse $\tau$, $T\sim 1/\tau$. Another approach Berges et al. (2018, 2018) is based on performing the trace over unobserved degrees of freedom in an expanding string described by the Schwinger model. In this case the effective temperature was found to be proportional to the inverse proper time $\tau$, $T=1/(2\pi\tau)$ Berges et al. (2018, 2018).
The maximal entanglement emerges from the evolution equations of high energy QCD at small Bjorken $x$ Kharzeev and Levin (2017); Kharzeev (2021). It manifests itself via a simple Boltzmann-like relation between the parton structure function $xG(x)$ and the entropy of the produced hadron system $S_{h}=-\sum p_{n}\ln p_{n}$ that is fixed by the independently measured multiplicity distribution $p_{n}$, $S_{h}=\ln xG(x)$ Kharzeev and Levin (2017). This relation was tested in experiment Tu et al. (2020); Andreev et al. (2021), and found to hold Kharzeev and Levin (2021); Hentschinski and Kutak (2022); Hentschinski et al. (2023, 2024). Maximal entanglement and conformal symmetry Lipatov (1994); Gürsoy et al. (2024) imply that the small $x$ behavior of the gluon structure function to be $xG(x)\sim x^{-1/3}$ in high-energy QCD Grieninger et al. (2025). The fragmentation of high-energy jets was also found to lead to the maximal entanglement Datta et al. (2025).
@@
A microscopic, first-principle, understanding of this emergent thermality within QCD is currently very difficult, as they require real-time Hamiltonian simulations in $(3+1)$ dimensions. Recent studies have therefore turned to lower-dimensional gauge theories where entanglement measures can be computed explicitly. Among these, the massive Schwinger model provides a minimal realization of confinement in $(1+1)$ dimensions, offering analytical tractability and access to large-scale tensor-network simulations. Indeed, the Schwinger model coupled to external sources was proposed as a way to describe the hadronization of jets by Casher, Kogut and Susskind long time ago Casher et al. (1974). Analytical solution of this model in the massless fermion limit was found in Kharzeev and Loshaj (2013).
Numerical real-time simulations of the massive Schwinger model coupled to external sources that describe the jets produced in $e^{+}e^{-}$ annihilation (using tensor-networks and exact-diagonalization) were recently performed in Florio et al. (2025, 2023, 2024); Janik et al. (2025); Barata and Rico (2025). These studies have shown that entanglement entropy (EE) grows rapidly during jet fragmentation. The physical content of the Schmidt states (eigenstates of the reduced density matrix) changes as a function of the coupling strength and proper time. While at early times and/or weak coupling the Schmidt eigenstates are close to the Fock quark-antiquark states, at late times and/or strong coupling they represent bound meson states Florio and Kharzeev (2021); Florio et al. (2025).
The crucially important finding of Florio et al. (2025) is that at late times the reduced density matrix of the system becomes very close to the thermal one. The resulting effective temperature appears universal, i.e. the comparison of the expectation values of all operators studied to the thermal ones yields the same value of the temperature Florio et al. (2025). These results indicate that entanglement not only underlies hadronization but may also drive the emergence of effective thermal behavior in confining gauge theories.
@@
Probing entanglement-induced thermality in stationary settings requires examining the entanglement structure of confining flux tubes. Previously, the tensor-network study Buyens et al. (2016) of the massive Schwinger model characterized confinement and string breaking through static observables and bipartite entanglement entropy, revealing clear signatures of flux tube formation and decay. On this basis, we test whether the reduced density matrix of states within the flux tube looks thermal, or whether the spatial entanglement can generate an effective thermality.
The process of string breaking has been extensively studied within both high-energy and condensed-matter physics. In spin-chain systems, recent theoretical works have shown that pairs of domain walls bound by a linear potential behave as analogs of particle-antiparticle strings, reproducing key signatures of flux tube formation and decay Verdel et al. (2020, 2023); Mallick et al. (2025); Artiaco et al. (2025). These concepts are now being tested experimentally in analog quantum simulators, where confinement dynamics have been observed in trapped-ion chains, Rydberg atom arrays, and superconducting qubit architectures Alexandrou et al. (2025); Crippa et al. (2024); Liu et al. (2025); De et al. (2024); Surace et al. (2024). Tensor-network and hybrid quantum–classical approaches have captured both static and dynamical aspects of string formation and breaking in one-dimensional $\mathbb{Z}_{2}$ gauge models Alexandrou et al. (2025), while recent extensions to two spatial dimensions demonstrate that confinement and screening persist in experimentally accessible geometries Borla et al. (2025); Cataldi et al. (2025); Xu et al. (2025); Di Marcantonio et al. (2025). Together, these developments establish a unified framework in which confinement and string breaking can be probed across high-energy, condensed-matter, and quantum-information platforms, setting the stage for a microscopic exploration of entanglement and thermality in static flux tubes.
@@ II The model
To analyze the quantum statistical properties of string breaking, we will use a model that shares many properties with QCD and is amenable to current quantum simulations – namely, the massive Schwinger model Schwinger (1962). The model is quite simple, but it exhibits the key features of confinement and chiral symmetry breaking.
@@
We study the process of string breaking by discretizing the theory on a spatial lattice, mapping the continuum theory onto a finite-dimensional Hilbert space. We adopt the Hamiltonian lattice formulation, introduced by Kogut and Susskind Kogut and Susskind (1975), in which the space is discretized while time remains continuous.
@@
[…]
where $\Theta$ is the Heaviside step function. In the absence of dynamical screening, this term introduces a region of a constant electric field between the two static charges, representing the confining flux tube connecting them.
To map the fermionic degrees of freedom to spin variables, we apply the Jordan-Wigner transformation Susskind (1977), which expresses the staggered fermion operators in terms of Pauli matrices:
@@ III Results
In the following, all observables are shown as functions of the dimensionless separation $d\cdot M_{s}$ between the two external static charges. Unless otherwise stated, all results correspond to $N=220,\,m=0.045\ a^{-1}$, and $g=0.09\ a^{-1}$. Throughout this work, we set $a=1$ and express temperatures and other dimensionful quantities in physical units defined by the mass of the first excited state of the Schwinger model, $M_{s}$, corresponding to the (pseudo)scalar meson. Our numerical results are obtained using the ITensor software library Fishman et al. (2022) in Julia Bezanson et al. (2015).
@@ III.2 Energy density and charge distribution
The spatial structure of the ground-state energy density, pressure, and charge density further illustrates the transition between the confined and screened regimes, as shown in Figure 3. In the ground state, the energy density and pressure are given by the expectation values of the $T^{00}$ and $T^{11}$ components of the energy momentum tensor, respectively (see for example Florio et al. (2025)).
@@ III.4 Entanglement entropy
The half-chain entropy, shown in the top panel of Figure 5, displays a striking non-monotonic behavior: it rises sharply with separation, reaches a peak near the critical distance, and then returns toward its vacuum value at large $d$. This peak provides a clear signature of confinement and its decay, consistent with the tensor-network results of Buyens et al. (2016), where the EE was identified as a sensitive probe of string formation and screening.
@@ III.6 Effective temperature from thermal overlaps
We quantify the degree of thermalization of the flux tube subsystem by comparing its reduced density matrix $\rho_{A}(d)$ to thermal density matrices $\rho_{\beta}$ of the static Hamiltonian at temperature $T=1/\beta$. Following the prescription introduced in Florio and Kharzeev (2021), we first obtain $\rho_{A}(d)$ from the ground state of the system with static sources separated by a distance $d$, tracing out all sites outside a centered region of length $L$. Separately, we construct thermal states for a system of length $N_{2}>L$ without external sources and compute their reduced thermal density matrices $\rho_{\beta}$ of a subsystem of the same size $L$. The comparison between $\rho_{A}$ and $\rho_{\beta}$ then determines the effective temperature $T(d)$.
@@
so that in both cases the effective temperature is identified by maximizing (8) and (10). Thermal wave functions are generated using the purification method (see Appendix D of Florio et al. (2025) for details).
@@
In both cases, the effective temperature increases with the separation between the external charges and reaches a pronounced maximum near the critical distance $d_{c}\cdot M_{s}\simeq 7$, where the string breaks. Note that as soon as the string breaks, the temperature reaches its vacuum value, even though the transition for the energy density is more smooth (presumably due to the residual interaction between the produced mesons). It is important to keep in mind that an effective temperature of the quantum vacuum is different from zero Florio et al. (2025).
@@ IV Discussion and outlook
Beyond the Schwinger model, the same mechanism is expected to operate in higher-dimensional confining gauge theories, including QCD. In those systems, the formation and subsequent breaking of flux tubes may similarly correspond to entanglement-driven thermalization processes. This correspondence offers a microscopic interpretation of the statistical features of hadronization and opens new avenues for exploring confinement dynamics through measures of quantum entanglement. In the future, it will be interesting to extend the study presented here to QCD; the lattice investigations of entanglement in confining flux tubes of Yang-Mills theory in $(1+1)$ and $(2+1)$ dimensions have already begun Amorosso et al. (2025, 2024).
@@ Acknowledgments
We are grateful to Adrien Florio, David Frenklakh and Shuzhe Shi for useful discussions and collaboration on related work. We also thank the participants of the QuantHep2025 workshop at Lawrence Berkeley National Lab for insightful comments. This work was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Grants No. DE-SC0020970 (Inqubator for Quantum Simulation (IQuS), S.G.), DE-FG02-97ER-41014 (UW Nuclear Theory, S.G.), DE-FG88ER41450 (SBU Nuclear Theory, D.K., E.M.) and by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under Contract No.DE-SC0012704 (S.G., D.K., E.M.). S.G. was supported in part by a Feodor Lynen Research fellowship of the Alexander von Humboldt foundation. E.M. was supported in part by the Center for Distributed Quantum Processing at Stony Brook University.
@@ Appendix A Symmetry resolved entanglement entropy
where $\lambda_{i}\geq 0$, $\sum_{i}\lambda_{i}=1$, and $\chi$ is the Schmidt rank. These coefficiente are the eigenvalues of both $\rho_{A}$ and $\rho_{B}$. From them, the EE follows as
@@
In the presence of a global $U(1)$ symmetry associated with fermion number conservation, the reduced density matrix of a subsystem commutes with the local charge operator $Q_{A}$, $\rho_{A},Q_{A}]=0$. As a result, $\rho_{A}$ is block-diagonal in the eigenvalues $Q_{A}$ of the conserved charge and can be decomposed as Turkeshi et al. (2020)

### v2

@@ I Introduction
Hadron production in high-energy collisions exhibits thermal features that appear universal and do not depend on the nature of colliding particles and the collision energy. The yields of different hadrons produced in $e^{+}e^{-}$, proton-proton, and heavy-ion collisions are well described by statistical models using a universal temperature [10, 5]. This apparent thermal behavior is very unlikely to originate from a conventional thermalization process involving rescattering of produced particles. For example, in $e^{+}e^{-}$ annihilation into hadrons, the produced hadrons (and the preceding quarks and gluons) are separated by a space-like interval and cannot undergo multiple final-state interactions.
Moreover, strong elliptical flow observed in heavy ion collisions at RHIC and LHC points towards a very fast thermalization – otherwise, the hydrodynamical evolution is too short and is not capable to reproduce the data, see [48] for a review. This early thermalization challenges the conventional models based on kinetic theory [48].
@@
Concrete realizations of this idea include the entanglement produced a rapid pulse of (chromo-)electric field during a high-energy collision [43, 42, 16, 27, 31, 30]. In this case an effective temperature $T$ appears proportional to the inverse duration of the pulse $\tau$, $T\sim 1/\tau$. Another approach [11, 11] is based on performing the trace over unobserved degrees of freedom in an expanding string described by the Schwinger model. In this case the effective temperature was found to be proportional to the inverse proper time $\tau$, $T=1/(2\pi\tau)$ [11, 11].
The maximal entanglement emerges from the evolution equations of high energy QCD at small Bjorken $x$ [38, 41]. It manifests itself via a simple Boltzmann-like relation between the parton structure function $xG(x)$ and the entropy of the produced hadron system $S_{h}=-\sum p_{n}\ln p_{n}$ that is fixed by the independently measured multiplicity distribution $p_{n}$, $S_{h}=\ln xG(x)$ [38]. This relation was tested in experiment [53, 4], and found to hold [39, 35, 33, 34]. Maximal entanglement and conformal symmetry [45, 32] imply that the small $x$ behavior of the gluon structure function to be $xG(x)\sim x^{-1/3}$ in high-energy QCD [29]. The fragmentation of high-energy jets was also found to lead to the maximal entanglement [20].
@@
A microscopic, first-principle, understanding of this emergent thermality within QCD is currently very difficult, as they require real-time Hamiltonian simulations in $(3+1)$ dimensions. Recent studies have therefore turned to lower-dimensional gauge theories where entanglement measures can be computed explicitly. Among these, the massive Schwinger model provides a minimal realization of confinement in $(1+1)$ dimensions, offering analytical tractability and access to large-scale tensor-network simulations. Indeed, the Schwinger model coupled to external sources was proposed as a way to describe the hadronization of jets by Casher, Kogut and Susskind long time ago [15]. Analytical solution of this model in the massless fermion limit was found in [40].
Numerical real-time simulations of the massive Schwinger model coupled to external sources that describe the jets produced in $e^{+}e^{-}$ annihilation (using tensor-networks and exact-diagonalization) were recently performed in [24, 26, 25, 36, 9]. These studies have shown that entanglement entropy (EE) grows rapidly during jet fragmentation. The physical content of the Schmidt states (eigenstates of the reduced density matrix) changes as a function of the coupling strength and proper time. While at early times and/or weak coupling the Schmidt eigenstates are close to the Fock quark-antiquark states, at late times and/or strong coupling they represent bound meson states [27, 24].
The crucially important finding of [24] is that at late times the reduced density matrix of the system becomes very close to the thermal one. The resulting effective temperature appears universal, i.e. the comparison of the expectation values of all operators studied to the thermal ones yields the same value of the temperature [24]. These results indicate that entanglement not only underlies hadronization but may also drive the emergence of effective thermal behavior in confining gauge theories.
@@
Probing entanglement-induced thermality in stationary settings requires examining the entanglement structure of confining flux tubes. Previously, the tensor-network study [14] of the massive Schwinger model characterized confinement and string breaking through static observables and bipartite entanglement entropy, revealing clear signatures of flux tube formation and decay. On this basis, we test whether the reduced density matrix of states within the flux tube looks thermal, or whether the spatial entanglement can generate an effective thermality.
The process of string breaking has been extensively studied within both high-energy and condensed-matter physics. In spin-chain systems, recent theoretical works have shown that pairs of domain walls bound by a linear potential behave as analogs of particle-antiparticle strings, reproducing key signatures of flux tube formation and decay [55, 56, 47, 6]. These concepts are now being tested experimentally in analog quantum simulators, where confinement dynamics have been observed in trapped-ion chains, Rydberg atom arrays, and superconducting qubit architectures [1, 19, 46, 21, 51]. Tensor-network and hybrid quantum–classical approaches have captured both static and dynamical aspects of string formation and breaking in one-dimensional $\mathbb{Z}_{2}$ gauge models [1], while recent extensions to two spatial dimensions demonstrate that confinement and screening persist in experimentally accessible geometries [18, 28, 13, 17, 57, 22]. Together, these developments establish a unified framework in which confinement and string breaking can be probed across high-energy, condensed-matter, and quantum-information platforms, setting the stage for a microscopic exploration of entanglement and thermality in static flux tubes.
@@
Note that the emergence of thermality and temperature is very natural in the holographic picture of a Schwinger pair [31, 30, 37, 50]. Since the fermion and anti-fermion are spatially separated, the minimal surface connecting the two (in the extra dimension) exhibits a world-sheet horizon which is associated with a temperature. Moreover, the thermal properties of bound states of the Schwinger model were discussed in [8, 7].
@@ II The model
To analyze the quantum statistical properties of string breaking, we will use a model that shares many properties with QCD and is amenable to current quantum simulations – namely, the massive Schwinger model [49]. The model is quite simple, but it exhibits the key features of confinement and chiral symmetry breaking.
@@
We study the process of string breaking by discretizing the theory on a spatial lattice, mapping the continuum theory onto a finite-dimensional Hilbert space. We adopt the Hamiltonian lattice formulation, introduced by Kogut and Susskind [44], in which the space is discretized while time remains continuous.
@@
$L_{\textrm{ext,}n}(d)=-\Theta\!\left(\frac{d/a-1}{2}-\left\lvert n-\frac{N}{2}\right\rvert+\delta\right)$ (5)
where $\Theta$ is the Heaviside step function and $\delta$ is a small number. In the absence of dynamical screening, this term introduces a region of a constant electric field between the two static charges, representing the confining flux tube connecting them.
To map the fermionic degrees of freedom to spin variables, we apply the Jordan-Wigner transformation [52], which expresses the staggered fermion operators in terms of Pauli matrices:
@@ III Results
In the following, all observables are shown as functions of the dimensionless separation $d\cdot M_{s}$ between the two external static charges. Unless otherwise stated, all results correspond to $N=220,\,m=0.045\ a^{-1}$, and $g=0.09\ a^{-1}$. Throughout this work, we set $a=1$ and express temperatures and other dimensionful quantities in physical units defined by the mass of the first excited state of the Schwinger model, $M_{s}$, corresponding to the (pseudo)scalar meson. Our numerical results are obtained using the ITensor software library [23] in Julia [12].
@@ III.2 Energy density and charge distribution
The spatial structure of the ground-state energy density, pressure, and charge density further illustrates the transition between the confined and screened regimes, as shown in Figure 3. In the ground state, the energy density and pressure are given by the expectation values of the $T^{00}$ and $T^{11}$ components of the energy momentum tensor, respectively (see for example [24]).
@@ III.4 Entanglement entropy
The half-chain entropy, shown in the top panel of Figure 5, displays a striking non-monotonic behavior: it rises sharply with separation, reaches a peak near the critical distance, and then returns toward its vacuum value at large $d$. This peak provides a clear signature of confinement and its decay, consistent with the tensor-network results of [14], where the EE was identified as a sensitive probe of string formation and screening.
@@ III.6 Effective temperature from thermal overlaps
We quantify the degree of thermalization of the flux tube subsystem by comparing its reduced density matrix $\rho_{A}(d)$ to thermal density matrices $\rho_{\beta}$ of the static Hamiltonian at temperature $T=1/\beta$. Following the prescription introduced in [27], we first obtain $\rho_{A}(d)$ from the ground state of the system with static sources separated by a distance $d$, tracing out all sites outside a centered region of length $L$. Separately, we construct thermal states for a system of length $N_{2}>L$ without external sources and compute their reduced thermal density matrices $\rho_{\beta}$ of a subsystem of the same size $L$. The comparison between $\rho_{A}$ and $\rho_{\beta}$ then determines the effective temperature $T(d)$.
@@
so that in both cases the effective temperature is identified by maximizing (8) and (10). Thermal wave functions are generated using the purification method (see Appendix D of [24] for details).
@@
In both cases, the effective temperature increases with the separation between the external charges and reaches a pronounced maximum near the critical distance $d_{c}\cdot M_{s}\simeq 7$, where the string breaks. Note that as soon as the string breaks, the temperature reaches its vacuum value, even though the transition for the energy density is more smooth (presumably due to the residual interaction between the produced mesons). It is important to keep in mind that an effective temperature of the quantum vacuum is different from zero [24].
@@ IV Discussion and outlook
Beyond the Schwinger model, the same mechanism is expected to operate in higher-dimensional confining gauge theories, including QCD. In those systems, the formation and subsequent breaking of flux tubes may similarly correspond to entanglement-driven thermalization processes. This correspondence offers a microscopic interpretation of the statistical features of hadronization and opens new avenues for exploring confinement dynamics through measures of quantum entanglement. In the future, it will be interesting to extend the study presented here to QCD; the lattice investigations of entanglement in confining flux tubes of Yang-Mills theory in $(1+1)$ and $(2+1)$ dimensions have already begun [3, 2].
@@ Acknowledgments
We are grateful to Adrien Florio, David Frenklakh and Shuzhe Shi for useful discussions and collaboration on related work. We also thank the participants of the QuantHep2025 workshop at Lawrence Berkeley National Lab for insightful comments. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Inqubator for Quantum Simulation (IQuS) under Award Number DOE (NP) Award DE-SC0020970 (S.G.). This work was also supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Grants No. DE-FG02-97ER-41014 (UW Nuclear Theory, S.G.), DE-FG88ER41450 (SBU Nuclear Theory, D.K., E.M.) and by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under Contract No.DE-SC0012704 (S.G., D.K., E.M.). S.G. was supported in part by a Feodor Lynen Research fellowship of the Alexander von Humboldt foundation. E.M. was supported in part by the Center for Distributed Quantum Processing at Stony Brook University. This work was also supported, in part, by the Department of Physics and the College of Arts and Sciences at the University of Washington.
@@ Appendix A Symmetry resolved entanglement entropy
where $\lambda_{i}\geq 0$, $\sum_{i}\lambda_{i}=1$, and $\chi$ is the Schmidt rank. These coefficients are the eigenvalues of both $\rho_{A}$ and $\rho_{B}$. From them, the EE follows as
@@
In the presence of a global $U(1)$ symmetry associated with fermion number conservation, the reduced density matrix of a subsystem commutes with the local charge operator $Q_{A}$, $\rho_{A},Q_{A}]=0$. As a result, $\rho_{A}$ is block-diagonal in the eigenvalues $Q_{A}$ of the conserved charge and can be decomposed as [54]